40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 17 00:24:47 2015, rana, jenne, Update, LSC, ALS Fool filter updated for more cancellation ALSfoo_150216.png15.png
    Reply  Wed Feb 18 01:58:51 2015, ericq, Update, LSC, ALS Fool single arm performance  ALSfool_spectra.pngALSfool_kick.pngALSfool_LoopShapes.pngALSfool_Feb182015.zip
       Reply  Wed Feb 18 17:51:40 2015, ericq, Update, LSC, ALS Fool impulse response ALSfool_fasterKick.png
          Reply  Wed Feb 18 19:06:40 2015, Koji, Update, LSC, ALS Fool impulse response latex-image-1.png
Message ID: 11041     Entry time: Tue Feb 17 00:24:47 2015     Reply to this: 11046
Author: rana, jenne 
Type: Update 
Category: LSC 
Subject: ALS Fool filter updated for more cancellation 

Today we measured the TFs again and then updated the filter in the POY -> ALS FF path so as to get 10x better cancellation.

The cancellation went from ~10 dB to ~30 dB. This seems good enough. The new filter 'Comp1' is just constructed by eye. We then had to tune the filter module gain to a few %. Seems good enough for now, but we should really try to understand what it is and why it is the way that it is. In the above plot, the ORANGE trace is the old cancellation and the GREEN one is the new one. The filter TF is attached below - its not special, we made it by presing buttons in FOTON until the TF matched the measured TF of ALSY/LSC-MC_CTRL_FF_OUT.

Attachment 1: ALSfoo_150216.png  173 kB  | Hide | Hide all
ALSfoo_150216.png
Attachment 2: 15.png  26 kB  | Hide | Hide all
15.png
ELOG V3.1.3-